OpenStudy (anonymous):

A company makes similar cylindrical containers in two sizes, as shown in the table below. Containers Radius (in centimeters) Volume (in cubic centimeters) Type 1 4 ? Type 2 7 1231.66 What is the volume, in cubic centimeters, of Type 1 containers? 2111.42 2155.41 402.17 229.81

OpenStudy (anonymous):

Please help! Pretty please?

OpenStudy (anonymous):

The answer I get is different than the answers provided, and I don't know what I am doing wrong.

OpenStudy (radar):

Cylinder container is like a circle with a skirt. To compute volume you would find the area of the circular base and multiply by the height. So lets look at type I fpr which I see a problem, I don't have the height. How tall are these cylinders?

OpenStudy (anonymous):

They don't provide the height.

OpenStudy (radar):

I think we are to use the relationship between the two types of cylinder assuming they are the same height, but different radius.

OpenStudy (anonymous):

Oh.... ok.

OpenStudy (anonymous):

What I tried to do was (pi x 4^2 / pi x 7^2) = (x/1231.66) But it didn't exactly work out, for the answer I got was 1226.29 = x. And that isn't labeled as one of the options.

OpenStudy (anonymous):

So..... It is quite a conundrum.

OpenStudy (anonymous):

So, may anybody please give me a clue, as what to do?

OpenStudy (radar):

So let us calculate the height of the Type 2. Volume=2pi r^2 h 1231.66 =2 pi 7 squared h\[1231.66 = 2\pi \times49\times h\]\[1231.66 \over 98 \pi = 4cm\] we can assume a height of 4 cm. Now for the volume of the type 1

OpenStudy (radar):

V(Type 1)=2 pi r^2 * 4 Volume would then equal:\[2 \pi \times 16 \times 4 =402.1234 cu. cm. \]

OpenStudy (anonymous):

OH... I never thought about it in that manner... Thank you... I, by myself would have never thought of finding the height through a reverse process. So then, by finding that out, you can plug it in to the formula to find out the answer we are seeking. It is more wise to view a problem through different views, so you could find all the different possibilities the problem in question may behold.

OpenStudy (radar):

My satellite link failed because of a thunderstorm that we are having. Sorry for delay.

OpenStudy (radar):

Did that value come close to one of your options?

OpenStudy (radar):

Hold on, we may be getting some new input from Callisto.

OpenStudy (anonymous):

Well, yes, it did come close to one of the answers.very close, in fact.

OpenStudy (radar):

I have a Pi key on calculator which makes it very accurate.

OpenStudy (anonymous):

Yes. The answer I have as one of the options provided, is rounded, so it came quite close. Only about a few hundredths off.

OpenStudy (radar):

@iluvmangos , remember this was based on an assumption that the cylinder containers were equal height. Hopefully the assumption was a valid one.

OpenStudy (anonymous):

Thank you. I am sure that it is. I will come on again in a few moments, and tell you if it is right or not.

OpenStudy (anonymous):

And unfortunately it is marked wrong... but at least we tried. That's what counts.

OpenStudy (radar):

@Callisto , did you have a comment?

OpenStudy (radar):

I was hoping Callisto could straighten me out, obviously I have gone astray here.

OpenStudy (callisto):

I got some different actually..

OpenStudy (callisto):

*something

OpenStudy (anonymous):

Well, this is certainly something I will ponder over..... and @Callisto , may you present your approach to the situation?

OpenStudy (radar):

Did I err on the calculations or on the assumption?

OpenStudy (callisto):

I use property of similar figures to do it.. (ratio of sides)^3 = ratio of volume \[(\frac{4}{7})^3 = \frac{V}{1231.66}\] Solve V.

OpenStudy (callisto):

@radar I saw the problem.... ''Volume=2pi r^2 h'' It should be volume= pi r^2 h

OpenStudy (radar):

Oh yes, I really boo booed! A=pi*r^2

OpenStudy (radar):

I think I combined the formula for circumfurence and area lol

OpenStudy (anonymous):

Oh goodness.... I guess that is where we went awry. Oops! ;)

OpenStudy (radar):

And as usual there was a distractor that was close to my incorrect answer. Sorry about that iluvmangos.

OpenStudy (radar):

Thanks Callisto for clarifying that for me. I will try not to do that any more.!!

OpenStudy (radar):

I like your method Callisto as it eliminates having to use Pi.

OpenStudy (anonymous):

Thanks for all the help you two have given me!

OpenStudy (radar):

Did you get around 230 cu. cm for volume and was that an option.

OpenStudy (radar):

You are welcome and sorry about the miscue.

OpenStudy (callisto):

The followings are quite useful: For similar figures, (ratio of sides)^2 = ratio of area (ratio of sides)^3 = ratio of volume Welcome and welcome and welcome :)